A cube of mass 10 kg is at rest on a horizontal surface when it is subjected to a horizontal force P that varies with time t as P = 20t. If the static and kinetic coefficients of friction are both 0.4, calculate the velocity of the cube after 5 s.

1
Expert's answer
2022-02-22T13:02:28-0500

"v=\int_0^t(P-\mu mg)dt=(10t^2-\mu mgt)|_0^t=50~\frac ms."